Solve the following system of linear equations by using the inverse matrix
method: (10)
x + y + z = 4
-2x - y + 3z = 1
y + 5z = 9
1
Expert's answer
2020-05-12T16:15:35-0400

Given system can be written in matrix form as "AX = b"

where "A = \begin{bmatrix}
1 & 1 & 1\\
-2 & -1 & 3 \\
0 & 1 & 5
\end{bmatrix}", "X = \begin{bmatrix}
x \\
y \\
z
\end{bmatrix}, 
b = \begin{bmatrix}
4 \\ 
1 \\ 
9
\end{bmatrix}".

Now, By expending with the first row "|A| = 1(-5-3)-1(-10)+1(-2) = -8+10-2 = 0"

So, "\rho(A)<3."

Now, "\begin{vmatrix}
1 & 1 \\
-2 & -1 
\end{vmatrix} = -1+2 = 1 \neq 0" .

So, "\rho(A) = 2" and number of variable to be let "= 3-\rho(A) = 1."

We let "z = k."

So, from given first two equations:

"x+y=4-k \\
-2x-y = 1-3k"

Which can be written as "\begin{bmatrix}
1 & 1\\
-2 & -1
\end{bmatrix} 
\begin{bmatrix}
x \\
y
\end{bmatrix} = 
\begin{bmatrix}
4-k\\
1-3k
\end{bmatrix}"

i.e. "A_1 X_1 = b_1" and "|A_1| = 1 \neq 0" .

So, Inverse method can apply in "A_1 X_1 = b_1",

"\implies X_1 = A_1^{-1} b_1"

"\implies 
\begin{bmatrix}
x \\
y
\end{bmatrix} = 
\begin{bmatrix}
-1 & -1\\
2 & 1
\end{bmatrix} 
\begin{bmatrix}
4-k\\
1-3k
\end{bmatrix}" "= \begin{bmatrix}
4k-5\\
-5k+9
\end{bmatrix}".

So, Solution of given system of equation is

"x =4k-5, y = -5k+9, z= k" where "k" is any real number.